# Deep Unfolding-Aided Parameter Tuning for Plug-and-Play Based Video Snapshot Compressive Imaging

### v1

@@ I INTRODUCTION
To take advantage of both methods, some hybrid approaches combining mathematical optimization and machine learning have been proposed for video SCI [27, 28, 29, 30, 31, 32, 33, 34, 35]. Most of these methods are based on the plug-and-play (PnP) approach [36, 37], where an update equation of the optimization algorithm is replaced with a denoising method for natural images or videos. For example, on the basis of generalized alternating projection (GAP) [38, 20] and alternating direction method of multipliers (ADMM) [39, 40, 41, 42], PnP-GAP and PnP-ADMM have been proposed for video SCI, respectively [27, 28]. For the incorporated denoising method, various techniques such as fast and flexible denoising network (FFDNet) [43] and fast deep video denoising net (FastDVDnet) [44] have been utilized. These PnP methods have lower computational complexity than that of the optimization-based DeSCI, while the learning-based high-quality denoising can be used in the algorithm.
@@
In this paper, we propose a deep unfolding-based tuning method for the noise level parameters in the PnP approach. Deep unfolding [45, 46, 47] leverages the similarity between the structure of iterative algorithms and neural networks to train appropriate values for internal algorithm parameters via automatic differentiation. Since the update equations of PnP-GAP and PnP-ADMM are differentiable with respect to the noise level parameters, deep unfolding can be used to train appropriate noise levels in principle. When we use deep learning-based denoising like FastDVDnet, however, we require huge GPU memory to train the noise level parameters. To reduce the required memory, we use the checkpointing technique [48, 49] in the proposed method. In computer simulations, we prepare the data for the training, train the noise level parameters, and evaluate the reconstruction performance with the conventional parameter settings. Simulation results show that the trained parameters via the proposed approach can achieve better reconstruction accuracy. Moreover, the trained parameters exhibit a non-monotonic behavior along the iterations of the reconstruction algorithms, which suggests that the monotonically decreasing noise levels assumed for the convergence proof [50, 28] are not optimal in terms of reconstruction accuracy. This interesting result shows similarity to the trained parameters of some other deep-unfolded algorithms [51, 52]. Our results would also provide insights into the research fields of deep unfolding.
@@ III-A PnP-ADMM
The first term in the problem (3) is a data fidelity term based on the measurement model in (2). The second term is a regularization term for videos and should be designed to return smaller values for clean videos. $\lambda$ ($>0$) is a parameter that balances the data fidelity term and the regularization term.
@@
Here, $\sigma_{k}$ is the noise level parameter to determine the strength of denoising. As the denoising function $\mathcal{D}_{\sigma_{k}}(\cdot)$, we can use various denoising methods such as FFDNet [43] and FastDVDnet [44]. In [28], FastDVDnet is used because it is trained for video denoising and can achieve high-quality denoising. Using powerful denoising methods improves the reconstruction accuracy of the algorithm empirically. To obtain good reconstruction accuracy, however, the parameters $\sigma_{k}\in[0,1]$ ($k=0,1,\ldots$) need to be appropriately chosen.
@@ III-B PnP-GAP
The constraint $\bm{y}=\bm{\Phi}\bm{x}$ restricts the variable $\bm{x}$ by using the measurement model (2). The variable $\bm{v}$ is optimized by minimizing the sum of the distance from $\bm{x}$ and the regularizer $R(\cdot)$ for videos. In PnP-GAP, we alternately update $\bm{x}$ and $\bm{v}$ in accordance with (18). The update equations of PnP-GAP are thus given by
@@ III-C Noise Level Paremeters in PnP Methods
In PnP-ADMM and PnP-GAP, we have parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ to control the strength of denoising at each iteration. The setting of these parameters significantly affects the accuracy of video reconstruction. In the previous study [28], the following three settings are considered for 60 iterations:
@@
    $\sigma_{0}=50/255,\sigma_{k+1}=0.97\sigma_{k}$ ($k=0,\dotsc,K-2$)
@@ III-D Other Related Work
$\bm{w}^{k+1}=\bm{\Psi}\left(\bm{y}-\bm{\Phi}\left(\bm{v}^{k}-\bm{u}^{k}\right)\right)$$\bm{x}^{k+1}=\bm{v}^{k}-\bm{u}^{k}+\bm{\Phi}^{\top}\bm{w}^{k+1}$$\bm{v}^{k+1}=\mathcal{D}_{\sigma_{k}}(\bm{x}^{k+1}+\bm{u}^{k})$$\bm{u}^{k+1}=\bm{u}^{k}+\bm{x}^{k+1}-\bm{v}^{k+1}$$\bm{w}^{k+1}$$\bm{x}^{k+1}$$\bm{v}^{k+1}$$\bm{u}^{k+1}$
@@ IV-B2 Reparametrization
where $p_{k}$ is the actual trained parameter. For comparison, the number of iterations in PnP-ADMM and PnP-GAP is set to 60. The initial value of $p_{k}$ in the training is set so that the noise level parameter $\sigma_{k}$ follows the first setting in Section III-C. This means
@@ IV-B3 Training
In the experiments, we use FastDVDnet [44] as the denoising method $\mathcal{D}_{\sigma_{k}}(\cdot)$. Since there are 60 FastDVDnet layers in the unfolded networks, we require huge GPU memory for the usual training. We thus use the checkpointing approach [48, 49] to reduce GPU memory usage. In regular backpropagation calculations, all intermediate results from the forward pass are stored for use in the backward pass. Checkpointing reduces memory usage by storing only selected intermediate results from the forward pass. For the non-stored results, recalculation is performed during the backward pass using the stored intermediate results. Consequently, this approach needs longer training time than the normal approach.
@@ V Simulation Results and Discussion
In this section, we evaluate the performance of the PnP-based methods with the proposed noise level parameter tuning via computer simulations. The simulation is conducted by using a computer with Intel Core i9-13900, 64GB memory, and NVIDIA RTX A4500. The simulation code is implemented in PyTorch 2.2.
@@ V-A Trained Noise Level Parameters and Average Reconstruction Accuracy
In the figure, ‘proposed’ represents the noise level parameters obtained by the proposed method. For comparison, we also show the noise level parameters in the first, second, and third setting in Section III-C as ‘conventional (1)’, ‘conventional (2)’, and ‘conventional (3)’, respectively. Note that the noise level parameters in the first setting are the initial values in the training. From Fig. 3, we can see that the noise level parameters change significantly by the training. The trained parameters is not necessarily monotonically decreasing, though the convergence of PnP-ADMM is usually guaranteed for a decreasing sequence of noise level parameters [50, 27, 28]. As shown in later, however, the proposed noise level parameters can achieve higher reconstruction accuracy than the conventional ones. This result suggests that monotonically decreasing noise level parameters are not necessarily optimal in terms of the final reconstruction accuracy. A similar phenomena to the zigzag behavior of the parameters has also been observed in the training of the TISTA [51], which is a deep unfolding-based method for sparse signal recovery.
@@ V-B Reconstruction Results for Each Video
A possible reason for the performance degradation of the conventional parameters is that the initial value of the algorithm is simply set as $\bm{v}^{0}=\bm{\Phi}^{\top}\bm{y}$ for the simple comparison. If we use the initizalization with GAP-total variation (GAP-TV) as in [28], the performance of the conventional parameters would be improved. By the proposed approch, however, we can obtain better reconstruction results without such initialization and reduce the computational cost. It should be noted that the proposed parameter tuning can also be combined with such initialization method.
@@
conventional (1) 27.15/0.832 20.01/0.685 30.07/0.798 32.89/0.817 24.21/0.800 26.60/0.791 26.82/0.787
conventional (2) 22.99/0.707 18.50/0.564 28.25/0.779 31.76/0.809 23.77/0.777 25.04/0.756 25.05/0.732
conventional (3) 22.11/0.716 17.39/0.497 24.11/0.734 16.88/0.272 17.55/0.437 20.49/0.647 19.76/0.550
proposed 30.84/0.894 27.16/0.879 32.91/0.889 36.42/0.918 27.65/0.887 28.13/0.858 30.52/0.887
@@
conventional (1) 24.26/0.530 19.06/0.498 23.84/0.414 24.73/0.408 22.03/0.436 23.04/0.439 22.83/0.454
conventional (2) 23.39/0.549 18.72/0.486 25.26/0.502 26.32/0.488 22.74/0.498 23.45/0.486 23.31/0.501
conventional (3) 19.19/0.405 15.23/0.326 20.08/0.351 15.60/0.169 15.74/0.260 16.11/0.312 16.99/0.304
proposed 30.04/0.833 26.45/0.829 31.23/0.803 33.52/0.816 27.15/0.817 27.66/0.788 29.34/0.814
@@
Moreover, Tables V and VI shows the performance of PnP-ADMM and PnP-GAP when the standard deviation of the Gaussian measurement noise is $0.02$, respectively.
TABLE V: PSNR (left entry) and SSIM (right entry) of PnP-ADMM. (noisy case, standard deviation: $0.02$)
@@
conventional (1) 23.30/0.615 17.99/0.470 25.51/0.615 27.99/0.636 23.04/0.665 24.74/0.653 23.76/0.609
conventional (2) 20.69/0.487 17.71/0.440 24.05/0.559 27.13/0.606 22.81/0.635 23.59/0.590 22.66/0.553
conventional (3) 21.06/0.535 16.78/0.411 22.95/0.559 16.74/0.246 17.34/0.368 20.26/0.552 19.19/0.445
proposed 28.22/0.782 24.53/0.746 29.28/0.738 31.33/0.749 26.68/0.791 27.16/0.766 27.87/0.762
TABLE VI: PSNR (left entry) and SSIM (right entry) of PnP-GAP. (noisy case, standard deviation: $0.02$)
@@
conventional (1) 21.76/0.428 17.56/0.382 22.43/0.370 23.73/0.371 21.29/0.392 22.16/0.394 21.49/0.389
conventional (2) 20.88/0.403 17.73/0.396 22.74/0.407 24.28/0.412 21.92/0.439 22.52/0.424 21.68/0.414
conventional (3) 18.75/0.343 14.80/0.282 19.27/0.300 15.38/0.156 15.55/0.232 15.98/0.291 16.62/0.267
proposed 27.39/0.712 23.85/0.702 27.92/0.650 29.16/0.649 26.00/0.696 26.46/0.671 26.80/0.680
From the tables, we can see that the proposed parameter tuning can achieve higher reconstruction accuracy than the conventional ones for most videos, even though the parameter is trained by using the data with the standard deviation of $0.01$. This result suggests that the proposed parameter tuning is robust to the change of the standard deviation of the measurement noise.
@@ V-C Evaluation for Real Data
In the figures, the top row shows the reconstruction results for each frame of conventional first parameter setting in Section III-C, and the bottom row shows the reconstruction results of the proposed parameters. In the experiments, $N_{t}=10$ frames of the video are reconstructed from a single compressed image. To obtain better reconstruction results, we use the initialization with GAP-TV for the real data. From Figs. 8 and 9, we can see that the proposed parameter tuning can achieve better reconstruction results than the conventional ones for real data, even though the parameter is trained by using the synthetic data. However, since there are still some artifacts in the reconstruction results (e.g., the latter frames in the domino video), further improvement would be required for the real data. One possible approach is to use the joint denoising strategy [23], where the weighted sum of multiple denoiser is use in the denoinsing process. Although we do not focus on the joint denoising strategy in this paper, the proposed parameter tuning can also be combined with the strategy.
@@ VI Conclusion
In this study, we have focused on the design of the noise level paremters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ in the PnP-based video SCI algorithms. We have proposed the data-driven tuning on the basis of the deep unfolding approach. In the proposed method, we prepare the dataset for the video SCI and train the parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ by minimizing the loss function. The simulation results show that the proposed noise level parameter can achieve higher reconstruction accuracy than the conventional ones. Moreover, since the parameters trained on noisy data can achieve good performance even on noiseless data, they demonstrate robustness to measurement noise. Since the performance improvement is obtained by the choice of the noise level parameters only, our results demonstrate the importance of the parameter tuning in PnP-based video SCI.

### v2

@@ I INTRODUCTION
To take advantage of both methods, some hybrid approaches combining mathematical optimization and machine learning have been proposed for video SCI [27, 28, 29, 30, 31, 32, 33, 34, 35]. Most of these methods are based on the plug-and-play (PnP) approach [36, 37], where an update equation of the optimization algorithm is replaced with a denoising method for natural images or videos. For example, on the basis of generalized alternating projection (GAP) [38, 20] and alternating direction method of multipliers (ADMM) [39, 40, 41, 42], PnP-GAP and PnP-ADMM have been proposed for video SCI, respectively [27, 28]. For the incorporated denoising method, various methods such as fast and flexible denoising network (FFDNet) [43] and fast deep video denoising net (FastDVDnet) [44] have been utilized. These PnP methods have lower computational complexity than that of the optimization-based DeSCI, while the learning-based high-quality denoising can be used in the algorithm.
@@
In this paper, we propose a deep unfolding-based tuning method for the noise level parameters in the PnP approach. Deep unfolding [45, 46, 47] leverages the similarity between the structure of iterative algorithms and neural networks to train appropriate values for internal algorithm parameters via automatic differentiation. Since the update equations of PnP-GAP and PnP-ADMM are differentiable with respect to the noise level parameters, deep unfolding can be used to train appropriate noise levels in principle. When we use deep learning-based denoising like FastDVDnet, however, we require huge GPU memory to train the noise level parameters. To reduce the required memory, we use the checkpointing technique [48, 49] in the proposed method. In computer simulations, we prepare the data for the training, train the noise level parameters, and evaluate the reconstruction performance with the conventional parameter settings. Simulation results show that the trained parameters via the proposed approach can achieve better reconstruction accuracy. Moreover, the trained parameters exhibit a non-monotonic behavior along the iterations of the reconstruction algorithms. This implies that the monotonically decreasing noise levels assumed for the convergence proof [50, 28, 34, 35] are not necessarily optimal in terms of reconstruction accuracy. This interesting result shows similarity to the trained parameters of some other deep-unfolded algorithms [51, 52]. Our results would also provide insights into the research fields of deep unfolding.
@@ III-A PnP-ADMM
The first term in the problem in (3) is a data fidelity term based on the measurement model in (2). The second term is a regularization term for videos and should be designed to return smaller values for clean videos. $\lambda$ ($>0$) is a parameter that balances the data fidelity term and the regularization term.
@@
Here, $\sigma_{k}\in[0,1]$ is the noise level parameter to determine the strength of denoising. As the denoising function $\mathcal{D}_{\sigma_{k}}(\cdot)$, we can use various denoising methods such as FFDNet [43] and FastDVDnet [44]. In [28], FastDVDnet is used because it is trained for video denoising and can achieve high-quality denoising. Using powerful denoising methods improves the reconstruction accuracy of the algorithm empirically. To obtain good reconstruction accuracy, however, the parameters $\sigma_{k}\in[0,1]$ ($k=0,1,\ldots$) need to be appropriately chosen.
@@ III-B PnP-GAP
The constraint $\bm{y}=\bm{\Phi}\bm{x}$ restricts the variable $\bm{x}$ by using the measurement model in (2). The variable $\bm{v}$ is optimized by minimizing the sum of the distance from $\bm{x}$ and the regularizer $R(\cdot)$ for videos. In PnP-GAP, we alternately update $\bm{x}$ and $\bm{v}$ in accordance with (18). The update equations of PnP-GAP are thus given by
@@ III-C Noise Level Paremeters in PnP Methods
In PnP-ADMM and PnP-GAP, we have parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ to control the strength of denoising at each iteration. The setting of these parameters significantly affects the accuracy of video reconstruction. Although FFDNet and FastDVDnet themselves can use noise level map with the same size as the input image, the noise level parameters in PnP methods are usually treated as scalar values, which means that each element of the noise level map is the same value among all pixels. In the previous study [28], the following three settings are considered for 60 iterations:
@@
    $\sigma_{0}=50/255,\sigma_{k}=0.97\sigma_{k-1}$ ($k=1,\dotsc,K-1$)
@@ III-D Other Related Work
$\bm{w}^{k+1}=\bm{y}-\bm{\Phi}\left(\bm{v}^{k}-\bm{u}^{k}\right)$$\bm{x}^{k+1}=\bm{v}^{k}-\bm{u}^{k}+\bm{\Phi}^{\top}\bm{\Psi}\bm{w}^{k+1}$$\bm{v}^{k+1}=\mathcal{D}_{\sigma_{k}}(\bm{x}^{k+1}+\bm{u}^{k})$$\bm{u}^{k+1}=\bm{u}^{k}+\bm{x}^{k+1}-\bm{v}^{k+1}$$\bm{w}^{k+1}$$\bm{x}^{k+1}$$\bm{v}^{k+1}$$\bm{u}^{k+1}$
@@ IV-B2 Reparametrization
where $p_{k}$ is the actual trained parameter. Note that $\sigma_{k}$ in (32) is in the range $(0,1)$. For comparison, the number of iterations in PnP-ADMM and PnP-GAP is set to 60. The initial value of $p_{k}$ in the training is set so that the noise level parameter $\sigma_{k}$ follows the first setting in Section III-C. This means
@@ IV-B3 Training
In the experiments, we use FastDVDnet [44] as the denoising method $\mathcal{D}_{\sigma_{k}}(\cdot)$. Since there are 60 FastDVDnet layers in the unfolded networks, we require huge GPU memory for the usual training. We thus use the checkpointing approach [48, 49] to reduce GPU memory usage, which can be implemented with a function of PyTorch. In regular backpropagation calculations, all intermediate results from the forward pass are stored for use in the backward pass. Checkpointing reduces memory usage by storing only selected intermediate results from the forward pass. For the non-stored results, recalculation is performed during the backward pass using the stored intermediate results. Consequently, this approach needs longer training time than the normal approach.
@@ V Simulation Results and Discussion
In this section, we evaluate the performance of the PnP-based methods with the proposed noise level parameter tuning via computer simulations. The aim of the simulation is to demonstrate the effect of the parameter settings, and hence we do not focus on the performance comparison with other SCI reconstruction methods. Note again that the proposed approach can be combined with some other approaches as discussed in Section III-D. The simulation is conducted by using a computer with Intel Core i9-13900, 64GB memory, and NVIDIA RTX A4500. The simulation code is implemented in PyTorch 2.2.
@@ V-A Trained Noise Level Parameters and Average Reconstruction Accuracy
In the figure, ‘proposed’ represents the noise level parameters obtained by the proposed method. For comparison, we also show the noise level parameters in the first, second, and third setting in Section III-C as ‘conventional (1)’, ‘conventional (2)’, and ‘conventional (3)’, respectively. Note that the noise level parameters in the first setting are the initial values in the training. In the simulations, we use the mask $\bm{M}$ in the dataset of the six grayscale videos. From Fig. 3, we can see that the noise level parameters change significantly by the training. The trained parameters is not necessarily monotonically decreasing, though the convergence of PnP-ADMM is usually guaranteed for a decreasing sequence of noise level parameters [50, 28, 34, 35]. As shown in later, however, the proposed noise level parameters can achieve higher reconstruction accuracy than the conventional ones. This result implies that monotonically decreasing noise level parameters might not be necessarily optimal in terms of the final reconstruction accuracy. A similar phenomena to the zigzag behavior of the parameters has also been observed in the training of the TISTA [51], which is a deep unfolding-based method for sparse signal recovery. Moreover, in the reinforcement learning-based parameter tuning for PnP methods [60], the resultant noise level parameters are not necessarily monotonically decreasing for medical resonance imaging (MRI).
@@ V-B Reconstruction Results for Each Video
A possible reason for the performance degradation of the conventional parameters is that the initial value of the algorithm is simply set as $\bm{v}^{0}=\bm{\Phi}^{\top}\bm{y}$ for the simple comparison. If we use the initizalization with GAP-total variation (GAP-TV) as in [28], the performance of the conventional parameters would be improved. By the proposed approch, however, we can obtain good reconstruction results without such initialization and reduce the computational cost. It should be noted that the proposed parameter tuning can also be combined with such initialization method.
@@
conventional (1) 29.83/0.924 22.73/0.857 34.24/0.946 41.02/0.981 25.16/0.894 27.40/0.880 30.06/0.914
conventional (2) 25.11/0.853 19.69/0.731 31.71/0.920 39.24/0.971 24.36/0.867 25.83/0.849 27.66/0.865
conventional (3) 22.40/0.790 17.60/0.570 24.26/0.849 16.90/0.283 17.63/0.474 20.50/0.693 19.88/0.610
proposed 32.32/0.938 28.43/0.930 35.45/0.949 41.51/0.986 28.04/0.924 28.54/0.887 32.38/0.936
@@
conventional (1) 25.88/0.608 21.03/0.619 25.75/0.511 26.20/0.472 23.00/0.513 23.77/0.491 24.27/0.536
conventional (2) 25.29/0.647 20.20/0.614 27.03/0.572 27.58/0.538 23.35/0.557 24.22/0.540 24.61/0.578
conventional (3) 19.42/0.457 15.37/0.365 20.40/0.444 15.69/0.182 15.82/0.282 16.22/0.353 17.15/0.347
proposed 31.90/0.914 28.01/0.912 34.36/0.922 37.72/0.921 27.75/0.890 27.84/0.831 31.26/0.898
@@
Moreover, TABLEs V and VI show the performance of PnP-ADMM and PnP-GAP when the standard deviation of the Gaussian measurement noise is $0.05$, respectively.
TABLE V: PSNR (left entry) and SSIM (right entry) of PnP-ADMM. (noisy case, standard deviation: $0.05$)
@@
conventional (1) 26.17/0.785 20.57/0.729 30.07/0.798 32.89/0.817 23.86/0.764 26.18/0.753 26.62/0.774
conventional (2) 22.14/0.644 18.74/0.600 28.25/0.779 31.76/0.809 23.48/0.739 24.68/0.717 24.84/0.715
conventional (3) 21.88/0.671 17.46/0.515 24.11/0.734 16.88/0.272 17.50/0.419 20.46/0.621 19.72/0.539
proposed 30.22/0.873 27.50/0.893 32.91/0.889 36.42/0.918 27.45/0.868 27.93/0.841 30.41/0.880
TABLE VI: PSNR (left entry) and SSIM (right entry) of PnP-GAP. (noisy case, standard deviation: $0.05$)
@@
conventional (1) 23.67/0.509 19.49/0.524 23.84/0.414 24.73/0.408 21.83/0.423 22.78/0.426 22.72/0.451
conventional (2) 22.64/0.512 18.99/0.512 25.26/0.502 26.32/0.488 22.56/0.485 23.17/0.468 23.16/0.494
conventional (3) 19.10/0.390 15.27/0.333 20.08/0.351 15.60/0.169 15.70/0.253 16.08/0.305 16.97/0.300
proposed 29.36/0.803 26.84/0.849 31.23/0.803 33.52/0.816 26.87/0.785 27.37/0.757 29.20/0.802
From the tables, we can see that the proposed parameter tuning can achieve higher reconstruction accuracy than the conventional ones for most videos, even though the parameter is trained by using the data with the standard deviation of $0.01$. This result suggests that the proposed parameter tuning is robust to the change of the standard deviation of the measurement noise. Alghough the performance in the noiseless case is slightly worse than that in the noisy case with $\sigma=0.01$ for some videos, this would be because the noise level parameters are trained by using the noisy data.
@@ V-C Evaluation for Real Data
In the figures, the top row shows the reconstruction results for each frame of conventional first parameter setting in Section III-C, and the bottom row shows the reconstruction results of the proposed parameters. In the experiments, $N_{t}=10$ frames of the video are reconstructed from a single compressed image. To obtain better reconstruction results, we use the initialization with GAP-TV for the real data. From Figs. 8 and 9, we can see that the proposed parameter tuning can achieve better reconstruction results than the conventional ones for real data, even though the parameter is trained by using the synthetic data. However, since there are still some artifacts in the reconstruction results (e.g., the latter frames in the domino video), further improvement would be required for the real data. One possible approach is to use the joint denoising strategy [23], where the weighted sum of multiple denoiser is used in the denoinsing process. Although we do not focus on the joint denoising strategy in this paper, the proposed parameter tuning can also be combined with the strategy.
@@ VI Conclusion
In this study, we have focused on the design of the noise level paremters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ in the PnP-based video SCI algorithms. We have proposed the data-driven tuning on the basis of the deep unfolding approach. In the proposed method, we prepare the dataset for the video SCI and train the parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ by minimizing the loss function. The simulation results show that the proposed noise level parameters can achieve higher reconstruction accuracy than the conventional ones. Moreover, the trained parameters can achieve good performance even when the standard deviation of the measurement noise is different from that in the training data. Since the performance improvement is obtained by the choice of the noise level parameters only, our results demonstrate the importance of the parameter tuning in PnP-based video SCI.